OpenStudy (anonymous):

prove (1-sin x)/sec x = cos^3 x/1+sin x any trig people in the houseeeeee

OpenStudy (ranga):

On LHS, replace 1/sec(x) with cos(x). Multiply top and bottom of LHS by (1 + sin(x)) and simplify to get the RHS.

OpenStudy (anonymous):

i have actually solved this working with both sides of the equation i just got lost in the steps a little bit so i was hoping someone could rework it

OpenStudy (anonymous):

=(12−sin2x)1cosx(1+sinx) =cos2x×cosx(1+sinx)=cos3x1+sinx

OpenStudy (anonymous):

how did we get this?

OpenStudy (ranga):

\[\frac{ 1-\sin(x) }{ \sec(x) } = (1 - \sin(x))\cos(x) = (1 - \sin(x))\cos(x) \times \frac{ (1+\sin(x)) }{ (1 + \sin(x)) } = ?\]

OpenStudy (anonymous):

i have this all typed up i am sorry it wont copy and paste right

OpenStudy (anonymous):

(1^2−sin^2x)/(1/cosx)(1+sinx)

OpenStudy (ranga):

(1 - sin(x)) * (1 + sin(x)) = 1^2 - (sin(x))^2 (because (a-b)(a+b) = a^2 - b^2 1^2 - (sin(x))^2 = cos^2(x)

OpenStudy (anonymous):

then multiply both sides buy cosx to eliminate that sec in the demoninator

OpenStudy (anonymous):

oh i think thats what i needed!

OpenStudy (ranga):

(1^2−sin^2x)/(1/cosx)(1+sinx) = cos^2(x) / { 1 / cos(x) * (1 + sin(x)) } = cos^3(x) / (1 + sin(x))

OpenStudy (anonymous):

i get it, its the algebra that kills me like that.

OpenStudy (ranga):

alright.

OpenStudy (anonymous):

great thanks!

OpenStudy (anonymous):

any last words of wisdom ill take from you haha

OpenStudy (ranga):

You are welcome. Dividing by a fraction is same as multiplying by its reciprocal. Here dividing by (1/cos(x)) is same as multiplying by cos(x).